OpenStudy (anonymous):

The height of a cylinder with a fixed radius of 10 cm is increasing at the rate of 0.5 cm/min. Find the rate of change of the volume of the cylinder (with respect to time) when the height is 30cm.

OpenStudy (anonymous):

dh/dt is given ... find dv/dt ... dv/dt = pi * r^2 *dh/dt r=10 cm ; dh/dt=o.5 cm/min

OpenStudy (phi):

you should start by writing down the formula for the volume of a cylinder

OpenStudy (phi):

\[ v= \pi r^2 h \] fixed radius of 10 cm means r is constant. height increases and volume increases. those variables are not constant. take the derivative with respect to time \[ \frac{d}{dt} v = \pi r^2\frac{d}{dt} h \] we treat pi and r as constants

OpenStudy (phi):

or simply \[ \frac{dv}{dt} = \pi r^2\frac{dh}{dt} \] or \[ \dot{v} = \pi r^2 \dot{h}\] (in physics, the dot over the variable is short-hand for derivative with respect to time. It was the notation Newton used) now replace the variables with the given values